On January 2, 2018, Jensen Corporation sells equipment it manufactured to Lewisburg Fabricators in exchange for a $200,000 note due in three years. The note bears no stated interest rate, but requires the entire $200,000 to be repaid at the end of three years. When Lewisburg Fabricators sought bank financing for this purchase the company was offered the funds at 8%, but decided to let Jensen hold the note.

A) What price does Jensen record the equipment sale at?

B) How much interest income is recorded in total the first two years (2018 and 2019)?

C) What is the balance of the note receivable two years later (end of 2019)?
Answer #1

Equipment sale will be recorded at the present value of payment to be received in future

= 200,000/(1.08)^3

= $158,766.45

B) Interest income in first year = 158,766.45*8% = $12,701.32

In Second year = (158766.45+12701.32)*8% = $13,717.42

C)Balance after two years = $158766.45+12701.32+13717.42 = $185,185.19